OpenStudy (anonymous):

Ok, 4x+1-5x=5-(x+4). Could someone help me with this one? I came up with x=-4 but the answer is "all real numbers" I don't see how they got that? All my odd-numbered problems are in the back of the book but they don't show you how to get to the solution:) Denise, Help!

OpenStudy (anonymous):

Solving the left side: 4x+1-5x = -x+1 Solving the right side: 5-(x+4) = 5-x-4 = 1-x which can be written as -x+1 Now lets compare: -x+1 = -x+1 Same equations on both side thus the answer is all real numbers.

OpenStudy (anonymous):

Ok, thanks for showing me how to get to the answer, I so appreciate it;) Denise